All three sides of the triangle shown are equal in length. Find the perimeter of the triangle. Side a is 2x + 5
Side b is 4x - 3
Side c is unknown

Answer by partha_ban(41)

Since side a = side b = side c,
4x - 3 = 2x + 5
4x - 2x = 5 + 3
2x = 8
x = 8 / 2 = 4
Therefore each side = 2*4 + 5 = 13
Hence, perimeter = 3a = 3*13 = 39